OpenStudy (anonymous):

Show all steps to solved the problem Be sure you check your solution for correctness If 3a-9=6 What is the value of 4a-1?

OpenStudy (kaederfds):

3a-9=6 3a=6+9 3a=15 a=5 4a-1=? 20-1=? =19